A 70kg volleyball player generates an average vertical ground reaction force of 1548N over a 0.47 second period of time while jumping to block an opposing player. What impulse was generated during this motion? How high will she jump?
Answer #1

here,

mass of volleyball , m = 70 kg

the force applied , F = 1548 N

time taken , t = 0.47 s

the impuse delivered , I = 1548 /0.47 kg.m/s

I = 727.56 kg.m/s

let the speed of player after the jump be v

I = m * v

727.56 = 70 * v

solving for v

v = 10.4 m/s

the speed of player after the jump is 10.4 m/s

the maximum height , hm = v^2 / 2g

hm = 10.4^2 /(2*9.81)

hm = 5.5 m